the segments formed by the altitude to the hypotenuse of right triangle ABC measures 8 inches and 10 inches. Find the length of the shorter leg of triangle ABC
Answer by dabanfield(803)

the segments formed by the altitude to the hypotenuse of right triangle ABC measures 8 inches and 10 inches. Find the length of the shorter leg of triangle ABC
Let x be the unknown side. By the Pythagoreon Theorem we have:
10^ = 8^2 + x^2
x^2 = 100 - 64 = 36
x = 6